Project

General

Profile

NDOS Commissioning Meeting 2012/04/02

Attendance: Jon, Mark Mathis, Minerba, Andrew Norman, Peter, Susan, Athans

Agenda:

  1. Beam and Detector Update
  2. DataCheck Report
  3. Installation Plans for the week
  4. DAQ Update and Plans for the week
  5. DCS Update and Plans for the week
  6. AOB

Beam and Detector Report

  • Jon got a run going Friday night and the DAQ ran fine through the weekend. A DSO scan was done this morning, and a new run started.

DataCheck Report:

  • no report this week

Installation plans for this week

  • no report

DAQ Update and Plans for next week

  • Peter: perhaps some more DDS testing, perhaps multiple partitions. Also trying to get simulated system running at FD, but is having problems.

DCS Update and Plans for next week

  • Athans: some issues last week involving a missing shared library caused DCS data to be lost over a couple of days. This was also corrupting DCMs. Stephen found and fixed the problem quickly after it was found.
  • Jon has finished loading the DCS data that was previously cached in the novadaq home directory into the production DB.

AOB

  • none